OpenStudy (anonymous):

*Will fan and medal!* Simplify: 5/11 + 4/9. Answer must be a simplified fraction.

OpenStudy (anonymous):



OpenStudy (anonymous):

got a calculator ? http://www.wolframalpha.com/input/?i=+5%2F11+%2B+4%2F9

OpenStudy (anonymous):

\[\huge \frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\] is how you add fractions